OpenStudy (anonymous):

evaluate the integral 18t^(1/3)ln(t)

OpenStudy (anonymous):

u=lnx and dv=18t^(1/3)

OpenStudy (zzr0ck3r):

yes dv = 18t^1/3)dt

OpenStudy (zzr0ck3r):

so u = d(ln(x))/dt and v = int(dv) u = 1/x and v = 9/2*x^4

OpenStudy (anonymous):

okay any need to figure out what du is?

OpenStudy (zzr0ck3r):

sorry du = 1/t dt

OpenStudy (zzr0ck3r):

replace any x with t..sorry

OpenStudy (anonymous):

no worries so how would we put that into the formula?

OpenStudy (zzr0ck3r):

\[uv - \int\limits_{}^{} v * du\]

OpenStudy (zzr0ck3r):

u = ln(t) v = (9/2) *t^4 du = (1/t) dt

OpenStudy (zzr0ck3r):

\[(\ln(t)*(9/2)t^4 - \int\limits_{}^{} (9/2) * x^4 * (1/x) dt\]

OpenStudy (zzr0ck3r):

\[\ln(t)*(9/2)t^4- (9/2)\int\limits_{}^{}(t^4/t) = \ln(t)*(9/2)t^4- (9/2)\int\limits_{}^{}(t^3) dt\]

OpenStudy (zzr0ck3r):

\[\ln(t)*(9/2)t^4- (9/2)(1/4)t^4 = \ln(t)*(9/2)t^4- (9/8)t^4 \]

OpenStudy (anonymous):

would that be ln(t)*-(9/6)t^4?

OpenStudy (anonymous):

oopes i meant ln(t)*(-9/6)t^4

OpenStudy (anonymous):

or is that the wrong way of structuring it

OpenStudy (zzr0ck3r):

sec phone

OpenStudy (anonymous):

ok :)

OpenStudy (zzr0ck3r):

i would factor out (9/2) t^4 and get (9/2)*t^4(ln(t)-(1/4))

OpenStudy (zzr0ck3r):

\[(9/2)*t^4*(\ln(t)-(1/4))\]

OpenStudy (zzr0ck3r):

wait this is wrong I forgot about the 18 sec sorry im doing lots of stuff

OpenStudy (anonymous):

hmm homeworks saying it isnt correct maybe..

OpenStudy (anonymous):

oh okay take your time!!

OpenStudy (zzr0ck3r):

v = (3/4)t^(4/3)

OpenStudy (zzr0ck3r):

so (27/2)*t^(4/3)*ln(t) - (27/2)int(t^(1/3)dt

OpenStudy (zzr0ck3r):

= (27/2)*t^(4/3)*ln(t) - (81*t^(4/3))/8 + c

OpenStudy (zzr0ck3r):

(27/8) t^(4/3)(4ln(t)-3)+c

OpenStudy (anonymous):

could that be simplified anymore?

OpenStudy (zzr0ck3r):

depends on what you call simple:)

OpenStudy (zzr0ck3r):

sec let me see if wolfram wil do this... http://www.wolframalpha.com/input/?i=integrate+18t%5E%281%2F3%29ln%28t%29 this is a great place to look at some integrals...notice how I just typed in what I wanted at the top

OpenStudy (zzr0ck3r):

hit the show steps button.

OpenStudy (anonymous):

okay ill try that. i was just asking because my hw was saying that the answer you have me was incorrect :/

OpenStudy (anonymous):

i found a form that worked thank you so much!! that site is very helpful!